OpenStudy (lacris):

Can anyone please help me? I will give you a shiny medal!! c:

OpenStudy (lacris):

I'm not sure what I'm doing wrong, I did the frequency over the total number but I'm not sure what I did wrong :(

OpenStudy (lacris):

0.06 was the only I got correct :(

OpenStudy (helpppp):

@agent0smith Show me your math skills and help them

OpenStudy (lacris):

So for the mens I did: 41+24+13+14+6 = 98 so that the sum of frequencies and did: A. 41/98 = relative frequency of 0.41836

OpenStudy (lacris):

B. 24/98 = 0.2448 C. 13/98 = 0.13265 D. 14/98 = 0.14285 E. 6/98 = 0.0612244 Women: (sum of frequencies): 18+40 + 13+13+18 = 102 A.18/102=0.1764705 B. 40/102 = 0.39215 C. 13/102 = 0.12745 D. 13/102 = 0.12745 E.18/102=0.1764705 Am I just rounding them incorrectly? Or did my math go wrong?

OpenStudy (agent0smith):

Why does the table show only two d.p. but say to round to three?

OpenStudy (lacris):

I'm not sure :( the program said for me to: (Round to three decimal places as​ needed.) :/ So I kind of don't know if I rounded correctly..

OpenStudy (lacris):

Was my math correct? Or did I round it wrong? Was it supposed to be up to only 3 significant figures?

OpenStudy (agent0smith):

Your math looks fine. Idk.

OpenStudy (lacris):

I added more values, and got most correct but now I'm confused on significant figures now :/ two sig figs: 0.01?

OpenStudy (lacris):

I correct them, but I don't know why they all aren't up to the same significant figures :c :/

OpenStudy (agent0smith):

B. 24/98 = 0.2448 is not correctly rounded in the table

OpenStudy (agent0smith):

Now it's right

OpenStudy (lacris):

oh, okay, thank you :D but do you know what they actually mean by significant figures? or do they want it all to have the same number placement?

OpenStudy (agent0smith):

Where does it mention sig figs??

OpenStudy (lacris):

my teacher told me to do that, but in the program, it only talks about rounding to the three decimal places.

OpenStudy (agent0smith):

Sig figs are different to d.p. Usually used for science classes. http://www.purplemath.com/modules/rounding2.htm

OpenStudy (lacris):

Ahh, I see, since this was data processing i didn't have to go into significant figures.. woopsies. Thank you so much! :D :)

OpenStudy (agent0smith):

Welcome.